# Enabling fundamental understanding of Nature with novel binning methods for 2D histograms

### 31 Mar 2026 v1

@@ 1 Introduction
Visualizing two-dimensional distribution samples is essential for reporting any modern scientific result. The main tool for doing that is the 2D histogram, constructed by dividing the sample space into regions (bins) and using color to show the number or density of samples in a given bin. Yet, despite the immense importance of 2D histograms, for a long time, data visualization practitioners have faced debilitating stagnation. Most tools (e.g. Matplotlib Hunter (2007), ROOT Brun and Rademakers (1997)) limit them to simple rectangular or, at best, hexagonal bins. Some domain-specific binning methods, such as HEALPix Górski et al. (2005), exist, but their applications are naturally limited.
@@
of Nature Binning, abbreviated as funbin.
@@
Matplotlib-compatible Python package; in Section 4 we
@@ 2 Conceptual foundation
”thrice great”, the Emerald Tablet Trismegistus (-100). In turn, one of the most famous quotes
@@
In medieval and Renaissance times, this principle was adopted to state the similarity between the Microcosm and Macrocosm, for example, in Paracelsus (1567). It is now our turn, not dissimilar to Paracelsus, to base our practices on near-inscrutable writing from centuries long gone. In the context of this work, the Macrocosm represents the objects under study in the world. The Microcosm is, in turn, the interior region of a scientific work, and most importantly, the plots. Therefore, the adaptation of Hermeto-Paracelsian insight to scientific practice is to find visual, conceptual, and thematic parallels between objects under study and the way one decides to visualize them.
If we examine the types of distribution one usually aims to visualize, we readily find almost none of them to look like squares. Nature rarely produces -symmetric shapes, instead favoring fuzzy blobs. Therefore, the widespread use of square or rectangular bins comes into pronounced conflict with the vast majority of the Macrocosm. One might argue that hexagonal bins improve on the situation by being closer to a circular shape (continuous circular symmetry) while still tiling the plane. However, in our view, replacing unnatural symmetry with a slightly more complex is merely a band-aid solution.
@@ 4 Examples
how binning features (Microcosm) enhance one’s fundamental understanding of the
@@
the scientific value of the plotted data as it breaks the malignant repetitiveness
associated with conventional periodic tilings, described in Section 2.
@@ 4.1 Penrose tilings
Penrose tilings are a set of aperiodic tilings described in Penrose (1979). The first of them chronologically, P1 tiling consists of five shapes: three kinds of pentagons (they look the same but have different matching rules, so are considered different tiles), a star, a diamond, and a ”boat”. The imagery is hard to miss: aside from a literal star, the boat is essentially a broken star, perhaps disrupted by tidal forces or collision; the pentagon is nothing but a star with no rays coming out from it, an extinct star, with multiplicity of similar pentagons showing distinct endpoints of stellar evolution: black hole, neutron star, white dwarf; finally, the diamond can be interpreted as referring to the crystalline structure at the cores of carbon-rich white dwarf. Driven by these considerations, in Figure 1 we plot data from HYG database D. as a binned Hertzsprung-Russell diagram, using Penrose P1 tiling as bins.
@@
merger. It is not surprising, then, that Roger Penrose himself has contributed greatly
@@
in observing binary black hole mergers through gravitational wave interferometry Abbott and others (2016).
@@
parameter estimation data release LVK Collaboration (2023). We notice that the two
@@
ratio equal to the golden ratio , shown in a golden dashed line.
@@
proton darts. Also, the darts and kites can be reinterpreted as symbolic representations of
particle jets with different substructure: larger intensity in the core (kite) or in the shell
(dart) of the jet. For these reasons, in Figure 3 we use LHC simulation data
from LHC Olympics BlackBox1 dataset Kasieczka et al. (2019) with Penrose P2 binning. We use
FastJet Cacciari et al. (2012) to analyze simulated jets and plot the 2D histogram of the invariant
mass of the main dijet and cylinder (-) distance between its two jets.
The intent behind the LHCO dataset is to provide a benchmark for anomaly detection
@@ 4.2 Aperiodic monotile
As explained in Section 4.1, aperiodicity itself is a highly desirable property for binnings used in scientific plots. Until recently, only two-tile aperiodic tiling, such as P2 and P3, have been known. But in 2023, a single shape that can tile the whole plane with no repeats, an aperiodic monotile, was discovered Smith et al. (2024). It is also called einstein, from ”ein Stein”, meaning ”one stone” in German. The addition of singularity to aperiodicity has profound implications for the Hermeto-Paracelsian reasoning: the tiling is built on a single principle, yet it never repeats itself. Thus, it provides a geometrical foundation for the scientific method, which is also a singular principle, yet it is supposed to facilitate an endless non-repeating procession of scientific activity.
@@
2798 pulsars from ATNF Pulsar Catalog Manchester et al. (2005) 222https://www.atnf.csiro.au/research/pulsar/psrcat/
@@
scientific insights, funbin is capable of life advice.
In another pulsar-related application, in Figure 5 we use a different aperiodic monotile from the same family, ”hat” (Tile), as a pixel for -ray data from Fermi LAT Atwood and others (2009). Namely, we have selected GeV photons in a radius around the direction of PSR J2032+4127 () recorded during the year 2025. The data show slowly varying diffuse emission and three bright sources, pulsars J2021+4026, J2021+3651, and J2032+4127. The bottom right corner shows the Fermi LAT point-spread function around GeV energy if it were shaped like the hat tile.
@@ 4.3 World map
In this section, we achieve an explicit Microcosm/Macrocosm convergence by bringing together two spheres: Earth’s political map provides binning for stellar mass distribution across the local Universe within 200 Mpc from Biteau (2021). For the former, we use the Natural Earth dataset Earth and volunteers to extract administrative boundaries of all countries at m precision. Since Earth’s oceans do not have administrative boundaries, they would constitute a single giant bin, which is inspiring theoretically, but leads to a poorly readable plot. Therefore, we subdivide the ocean with the regular HEALPix bins (). For the data we extract from the Biteau (2021) galaxies with Mpc and weight them according to their estimated , accounting for the mass incompleteness correction. We then map their Galactic longitude and latitude to Earth’s longitude and latitude and juxtapose them on our binning. The resulting 2D histogram in Mollweide projection is shown in Figure 6.
@@
of Earth and Universe, we essentially rediscover Kantian epistemology: the Earth’s map serves
as a region of transcendental, a frame structuring our experience of the Cosmos. Perhaps there is
@@ 5 Discussion
portion of the problems (while also introducing interesting new ones)
@@ Acknowledgements.
- Philosophia magna, tractus aliquot, 1567. Paracelsus. Cited by: §2.

### 01 Apr 2026 v2

@@ 1 Introduction
Visualizing two-dimensional distribution samples is essential for reporting any modern scientific result. The main tool for doing that is the 2D histogram, constructed by dividing the sample space into regions (bins) and using color to show the number or density of samples in a given bin. Yet, despite the immense importance of 2D histograms, for a long time, data visualization practitioners have faced debilitating stagnation. Most tools (e.g., Matplotlib (Hunter, 2007), ROOT (Brun and Rademakers, 1997)) limit them to simple rectangular or, at best, hexagonal bins. Some domain-specific binning methods, such as HEALPix (Górski et al., 2005), exist, but their applications are naturally limited.
@@
of Nature Binning, abbreviated funbin.
@@
a Matplotlib-compatible Python package; in Section 4 we
@@ 2 Conceptual foundation
”thrice great”, the Emerald Tablet (Trismegistus, -100). In turn, one of the most famous quotes
@@
In medieval and Renaissance times, this principle was adopted to state the similarity between the Microcosm and Macrocosm, for example, in Paracelsus (1567). It is now our turn, not dissimilar to Paracelsus, to base our practices on near-inscrutable writing from centuries long gone. In the context of this work, the Macrocosm represents the objects under study in the world. The Microcosm is, in turn, the interior region of a scientific work, most importantly, the plots. Therefore, the adaptation of Hermeto-Paracelsian insight to scientific practice is to find visual, conceptual, and thematic parallels between objects under study and the way one visualizes them.
If we examine the types of distribution one usually aims to visualize, we readily find almost none of them to look like squares. Nature rarely produces -symmetric shapes, instead favoring fuzzy blobs. Therefore, the widespread use of square or rectangular bins comes into pronounced conflict with the vast majority of the Macrocosm. One might argue that hexagonal bins improve on the situation by being closer to a circular shape (continuous circular symmetry) while still tiling the plane. However, in our view, replacing the unnatural symmetry with a slightly less unnatural is merely a band-aid solution.
@@ 4 Examples
how the chosen binning (Microcosm) enhance one’s fundamental understanding of the
@@
the scientific value of the plotted data, since it breaks the malignant repetitiveness
associated with conventional periodic tilings, as described in Section 2.
@@ 4.1 Penrose tilings
Penrose tilings are a set of aperiodic tilings described in Penrose (1979). The first of them chronologically, P1 tiling consists of five shapes: three kinds of pentagons (they look the same but have different matching rules, so are considered different tiles), a star, a diamond, and a boat. The imagery is hard to miss: aside from a literal star, the boat is essentially a broken star, perhaps disrupted by tidal forces or collision; the pentagon is nothing but a star with no rays coming out from it, an extinct star, and the three similarly-looking pentagons are three different endpoints of stellar evolution: black hole, neutron star, white dwarf; finally, the diamond can be interpreted as referring to the crystalline structure at the cores of carbon-rich white dwarfs. Driven by these considerations, in Figure 1 we plot data on observed star color and magnitude from the HYG database (Nash, ) as a binned Hertzsprung-Russell diagram, using Penrose P1 tiling as bins.
@@
merger. It is not surprising then that Roger Penrose himself has contributed greatly
@@
in observing binary black hole mergers through gravitational wave interferometry (Abbott and others, 2016).
@@
parameter estimation data release (LVK Collaboration, 2023). We notice that the two
@@
ratio close to the golden ratio , shown in a golden dashed line.
@@
proton darts. Also, the dart and kite shapes can be reinterpreted as symbolic representations of
particle jets inside of the collider. Larger intensity in the central region of the jet
gives the kite-shaped jet profile, while larger intensity in the shell — the dart-like
one. For these reasons, in Figure 3 we plot LHC simulation data
from LHC Olympics 2020 BlackBox1 dataset (Kasieczka et al., 2019) with Penrose P2 binning. We use
FastJet (Cacciari et al., 2012) to analyze simulated jets and plot the total invariant mass
of the two heaviest jets against the - cylinder distance between them.
The central purpose of the LHCO dataset is to provide a benchmark for anomaly detection
@@ 4.2 Aperiodic monotile
As explained in Section 4.1, aperiodicity itself is a highly desirable property for binnings used in scientific plots. Until recently, only two-tile aperiodic tiling, such as P2 and P3, have been known. But in 2023, a single shape that can tile the whole plane with no repeats, an aperiodic monotile, was discovered (Smith et al., 2024). It is also called einstein, from ”ein Stein”, meaning ”one stone” in German. The addition of singularity to aperiodicity has profound implications for the Hermeto-Paracelsian reasoning: the tiling is built on a single principle (shape), yet it never repeats itself. Thus, it provides a geometrical foundation for the scientific method, which is also a singular principle, yet it is supposed to facilitate an endless non-repeating procession of scientific activity.
@@
2798 pulsars from ATNF Pulsar Catalog (Manchester et al., 2005) 222https://www.atnf.csiro.au/research/pulsar/psrcat/
@@
scientific insights, funbin is capable of giving life advice.
In another pulsar-related application, in Figure 5 we use a different aperiodic monotile from the same family, ”hat” (Tile), as a pixel for -ray data from Fermi LAT (Atwood and others, 2009). Namely, we have selected GeV photons in a radius around the direction of PSR J2032+4127 () recorded during the year 2025. The data show slowly varying diffuse emission and three bright sources, pulsars J2021+4026, J2021+3651, and J2032+4127. The bottom right corner shows the Fermi LAT point-spread function around GeV energy if it were shaped like the hat tile.
@@ 4.3 World map
In this section, we achieve an explicit Microcosm/Macrocosm convergence by bringing together two spheres: Earth’s political map provides binning for stellar mass distribution across the local Universe within 200 Mpc from Biteau (2021). For the former, we use the Natural Earth dataset (Natural Earth and others, ) to extract administrative boundaries of all countries at m precision. Since Earth’s oceans do not have administrative boundaries, they would constitute a single giant bin, which leads to a poorly readable plot. Therefore, we subdivide the ocean with the regular HEALPix bins (). For the data we extract from Biteau (2021) galaxies with Mpc and weight them according to their estimated , accounting for the mass incompleteness correction. We then juxtapose their Galactic longitude and latitude onto Earth’s longitude and latitude and our combined countires and ocean pixels binning. The resulting 2D histogram in the Mollweide projection is shown in Figure 6.
@@
of Earth and the Universe, we essentially rediscover Kantian epistemology: the Earth’s map serves
as a region of transcendental, a frame that structures our experience of the Cosmos. Perhaps there is
@@ 5 Discussion
portion of the problems (while also introducing interesting new ones),
@@ Acknowledgements.
- Philosophia magna, tractus aliquot. Paracelsus. Cited by: §2.